How much water needs to be mixed with a 15% dye solution to get 24 gallons of a solution that is 5% dye?
Answer by checkley77(12844)

.15(24-X)=.05*24
3.6-.15X=1.2
-.15X=1.2-3.6
-.15X=-2.4
X=-2.4/-.15
X=16 GALLONS OF 15% DYE IS USED.
24-16=8 GALLONS OF WATER IS USED.
PROOF:
.15(24-16)=.05*24
.15*8=1.2
1.2=1.2